OpenStudy (anonymous):

Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form

OpenStudy (anonymous):

find slope of y=-x+6 then because these lines are parallel , the other line will also have same slope use y-y1=m(x-x1) to find desired equation

OpenStudy (anonymous):

its a timed test help please :0

OpenStudy (anonymous):

Parallel lines have the same slope and you have a point. So you can use the point-slope form and change it into slope-intercept. (–4, 2) y - 2 = 1(x -(-4)) y - 2 = 1x + 4 y = 1x + 6 Hope this helps! (:

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

Your welcome! Practice makes Perfect...the more you do these kinds of problems on your own, the easier they will be(: